OpenStudy (anonymous):

Solve Each Equation by first clearing the fractions or decimals 1/3+1/9(k+5)-k/4=2 I Need to show all work

OpenStudy (anonymous):

\[\frac{ 1 }{ 3}+\frac{ 1 }{ 9 }(k+5)-\frac{ k }{4 }\]

OpenStudy (mertsj):

Can you identify the common denominator?

OpenStudy (anonymous):

y=1/3-3+1

OpenStudy (anonymous):

@Mertsj the common denominator is 36

OpenStudy (mertsj):

Yes, very good. So multiply each term by 36.

OpenStudy (mertsj):

\[\frac{1}{3}\times 36+\frac{k+5}{9}\times 36-\frac{k}{4}\times36=2\times36\]

OpenStudy (mertsj):

\[12+4(k+5)-9k=72\]

OpenStudy (mertsj):

Can you finish it?

OpenStudy (anonymous):

@Mertsj yes thank you

OpenStudy (mertsj):

yw